# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

### v2

@@ Appendix E Detailed Discussion on Countermeasures
Rowhammer defenses. Popular Rowhammer mitigation strategies include hardware-level defenses that employ row-refreshing and swapping [82, 83, 84, 16, 85] or error correcting codes (ECC) [86, 87]. However, row-refreshing and swapping incur significant area and latency costs, and ECC are bypassable with advanced Rowhammer attacks [62, 69, 41]. In contrast, system-level defenses modify memory allocation protocols [88, 89, 90, 70], but require changes to memory allocators and specialized memory layouts, which introduce substantial memory and computational overhead. The limitations and overhead have restricted their adoption in real-world infrastructure.

### v3

@@ Appendix G Detailed Discussion on Countermeasures
Rowhammer defenses. Rowhammer mitigation strategies include hardware-level defenses that employ row-refreshing and swapping [54, 74, 39, 42, 91] or error correcting codes (ECC) [20, 73]. However, row-refreshing and swapping incur significant area and latency costs, and ECC are bypassable with advanced attacks [18, 41, 31]. In contrast, system-level defenses modify memory allocation protocols [9, 5, 44, 83], but require changes to memory allocators and specialized memory layouts, which introduce substantial memory and computational overhead. The limitations and overhead have restricted their adoption in real-world infrastructure.